OpenStudy (anonymous):

if a and b are independent events then it must be true that P(A|B)=P(A) TRUE OR FALE

OpenStudy (juanpablojr):

What do you think?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

yeah...see events A and B are independent if any one of the following three equivalent conditions hold: P(A ∩ B) = P(A)P(B) P(A|B) = P(A) B has no effect on A P(B|A) = P(B) A has no effect on B Intuitively, two events are independent if the occurrence of one has no effect at all on the probability of the other. If two events A and B are not independent, then they are dependent....!!

OpenStudy (anonymous):

so true is right?

OpenStudy (anonymous):

yes,absolutely.

OpenStudy (anonymous):

woo thnsxs ur good at this

OpenStudy (anonymous):

hahah AM I? :)

OpenStudy (anonymous):

can u help me with this @aryandecoolest

OpenStudy (anonymous):

tomorrow for sure..i should leave now..it's 2 AM here..need to sleep :)